OpenStudy (anonymous):

limits as x approaches 0 (2+x)^3 -8/x

myininaya (myininaya):

What happens when you combine the fractions.

OpenStudy (anonymous):

it doesnt work

OpenStudy (accessdenied):

I think this is meant to be written as: \( \displaystyle \lim_{x \to 0} \left( \frac{(2+x)^3 - 8}{x} \right) \) The trick here is to recognize the difference of cubes in the numerator. We can factor things of that form like so: \( a^3 - b^3 = (a - b)(a^2 + ab + b^2) \) The first cube is (2+x)^3. The second cube you do not directly see, but recognize that 8 = 2^3 by prime factorization! \( \displaystyle \lim_{x \to 0} \left( \frac{(2+x)^3 - 2^3}{x} \right) \) By factoring the difference of cubes, it becomes: \( \displaystyle \lim_{x \to 0} \left( \frac{\color{red}{(\mathbf{2+x - 2})}((2+x)^2 + (2+x) \cdot 2 + 2^2)}{\mathbf{\color{red}{x}}} \right) \) Notice that the two highlighted factors will become the same, and cancel each other. The remaining factor can be solved by substitution!